OpenStudy (anonymous):

Evaluate f(x) = 2x5 -3x4 + x3 - 6x + 20: x = 2 1. 32 2. 10 3. 8 4. 4

myininaya (myininaya):

Did you try replacing x with 2?

myininaya (myininaya):

Then use order of operations

OpenStudy (anonymous):

ok super duper easy substitute 2 for x f(2)=2(2)^5-3(2)^4+2^3-6(2)+20 f(2)= ? Im pretty sure you can figure it ou from there

OpenStudy (anonymous):

okay i got 10, is that correct?

myininaya (myininaya):

I think you should try again I'm getting something different

OpenStudy (anonymous):

pkay hold on im horrible at math

OpenStudy (anonymous):

38?

myininaya (myininaya):

Lets try this by hand We have \[2(2)^5-3(2)^4+(2)^3-6(2)+20\] Do you know the order of operations?

myininaya (myininaya):

We have no grouping symbols But we do have exponents Evaluate all that has exponent first What do we get?

myininaya (myininaya):

\[2^5=?\] \[2^4=?\] \[2^3=?\]

OpenStudy (anonymous):

so do the exponents matter?

myininaya (myininaya):

Yes they matter

myininaya (myininaya):

That is why we are doing something with them

myininaya (myininaya):

\[2 \neq 2^5\]

myininaya (myininaya):

b/c \[2^5=32 \neq 2\] Do exponents matter big time

myininaya (myininaya):

Order of operations tells us to take care of the exponents so that is what we are doing I asked you 3 questions above can you answer them?

OpenStudy (theviper):

First put the value of x in the equation.Then what will u got is the ans.:)

OpenStudy (theviper):

\[\Huge{\color{red}{f(x)=2x^5-3x^4+x^3-6x+20}}\]\[\LARGE{\color{green}{f(2)=2(2)^5-3(2)^4+(2)^3-6(2)+20}}\]\[\LARGE{\color{green}{\implies(2\times32)-(3\times16)+8-12+20}}\]\[\LARGE{\color{green}{\implies64-48+8-12+20}}\]\[\Huge{\color{gold}{\star \star}{32Ans.}}\]